OpenStudy (anonymous):

Use Lagrange multipliers to find the maximum and minimum values of f(x,y)=2x+y, subject to the constraint x^2+y^2≤2

OpenStudy (anonymous):

fx = 2 fy = 1 gradF = <2,1> gradG = <2x,2y> 2 = L2x 1 = L2y 2y = x x^2 + 4x^2 = 2 5x^2 =2 x= +-sqrt(2/5) y = +-2*(sqrt(2/5)) max = 4*sqrt(2/5) min = -4*sqrt(2/5) wrong?

OpenStudy (anonymous):

@experimentX